OpenStudy (erinweeks):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 2 + 5 + 8 + . . . + ( 3n - 1) = n(1 + 3n)/2

ganeshie8 (ganeshie8):

u knw what S1 means ?

ganeshie8 (ganeshie8):

S1 means this :- sum of first 1 terms

ganeshie8 (ganeshie8):

S2 means, sum of first 2 terms S3 means, sum of first 3 terms ......... S1000 means, sum of first 1000 terms

ganeshie8 (ganeshie8):

we're given the series formula :- Sn = 2 + 5 + 8 + ....... + (3n+1) = n(1+3n)/2

ganeshie8 (ganeshie8):

use this, and can u calculate S1 ?

ganeshie8 (ganeshie8):

Sn = n(1+3n)/2 S1 = 1(1+3*1)/2 = ?

OpenStudy (erinweeks):

well S1 = 2?

OpenStudy (erinweeks):

sorry @ganeshie8 had to grab some to munch on lol

ganeshie8 (ganeshie8):

yup ! thats the sum of first 1 terms

ganeshie8 (ganeshie8):

ohk... np :)

ganeshie8 (ganeshie8):

now find sum of first 1 terms manually

OpenStudy (erinweeks):

what you mean by that

ganeshie8 (ganeshie8):

first term is 2 itself, so yes the Sn formula is correct for S1

ganeshie8 (ganeshie8):

2 + 5 + 8 + ....... + (3n+1) ^

ganeshie8 (ganeshie8):

Next, find S2 using the formula

ganeshie8 (ganeshie8):

Sn = n(1+3n)/2 S2 = ?

OpenStudy (erinweeks):

what do i plug in for n?

ganeshie8 (ganeshie8):

2

OpenStudy (erinweeks):

so S2 = 7

ganeshie8 (ganeshie8):

yup, now add the first two numbers in given sequence manually and see wat u get

ganeshie8 (ganeshie8):

2 + 5 + 8 + ....... + (3n+1) ^ ^

OpenStudy (erinweeks):

7!

ganeshie8 (ganeshie8):

so... S2 is true

ganeshie8 (ganeshie8):

next, find S3 using the formula

ganeshie8 (ganeshie8):

Sn = n(1+3n)/2 S3 = ?

OpenStudy (erinweeks):

15

ganeshie8 (ganeshie8):

now as usual, add the first 3 terms manually and see wat u get

OpenStudy (erinweeks):

15

ganeshie8 (ganeshie8):

So its same, so the formula is verified for S1, S2, S3

OpenStudy (erinweeks):

okay got it (:

ganeshie8 (ganeshie8):

cool :) also observe the general term :- Sn = 2 + 5 + 8 + ....... + (3n+1) = n(1+3n)/2 ^^^^^ we could have found the first few terms using that general term as well :- general term an = 3n-1 so, first term a1 = 3*1-1 = 2 second term a1 = 3*2-1 = 5 third term a3 = 3*3-1 = 8 ...

OpenStudy (erinweeks):

oh okay gotchaa (: